geerky42 (geerky42):

\[\Large \int\limits_{0}^{8}\int\limits_{0}^{y^{1/3}}\sqrt{x^4+1}dxdy\]

geerky42 (geerky42):

I changed the order of integration, then when I tried to evaluate it, I am stuck at \(\Large \displaystyle \int_{0}^2 8\sqrt{x^4+1} - x^3\sqrt{x^4+1} dx\) How can I solve that?

geerky42 (geerky42):

Any idea? @SithsAndGiggles

OpenStudy (anonymous):

Not for the first term. There's doesn't seem to be a closed form for it. The second term could be integrated with a substitution though.

OpenStudy (anonymous):

Have you tried converting to polar coordinates?

geerky42 (geerky42):

No, I haven't because I don't see how that would helps me.

geerky42 (geerky42):

Maybe I made mistake when I switch integral order, but I checked and rechecked, nothing is wrong. But it seem like it cannot be solved by hand.

OpenStudy (anonymous):

No, your work is right. And you're right about not being able to solve it by hand, though you can try to approximate it with the first few terms of a the power series for \(\sqrt{x^4+1}\). I don't think there's any easy way to integrate what you have so far. And you can ignore the polar-conversion suggestion, it only serves to make things I harder from the looks of it.

geerky42 (geerky42):

hmm, ok. thanks for helps though.